# When Worse is Better: Navigating the compression-generation tradeoff in visual tokenization

### v1

@@ 4 Analysis
Factors affecting validation loss. We saw that changing the number of tokens per image (Section 3.4) affected compute scaling more significantly than changing the codebook size (Section 3.5). Why is that? We note that for a unigram model, the loss lower bound is the empirical entropy of the codebook distribution per-position, since cross-entropy can be decomposed into $H(X)+D_{KL}(q(X)|p(X))$ and $D_{KL}\to 0$ when $q\to p$. In Figure 10 (top left), we measure this directly over the ImageNet train set over various codebook sizes. [^3] Note the periodicity, implying that positions near the center of the image have higher entropy. In Figure 10 (top right), we see that the average entropy per position converges, even as the overall codebook entropy increases log-linearly (as one would expect with 100% codebook utilization), implying that as codebook size increases, codes become more specialized by position. This has the secondary effect of ensuring that the lower bound for a unigram model over these tokens also does not increase, meaning that our stage 2 model can still learn even with very large codebook sizes.
Figure 10: Entropy analysis of image tokenizers with varying codebook sizes. (top left) Per-position token entropy following raster scan order, showing higher entropy at image centers and lower at edges. (top right) Global token distribution approaches uniform (diagonal), while per-position expected entropy shows systematic deviation. (bottom left) Distribution skew metric ($1-\frac{2^{\text{per position entropy}}}{2^{\text{total entropy}}}$) reveals optimal global codebook usage but convergent local usage as codebook size increases, allowing effective stage 2 causal model training. (bottom right) Effects of CRT tokenization uniformly reduce stage 2 loss across all positions. See Section 4 for more details.
What does CRT do? The lower-bound for an $n$-gram model is upper-bounded by this estimate, as $H(X_{i})\geq H(X_{i}|X_{0}\ldots X_{i-1})$. $H(X_{i}|X_{0}\ldots X_{i-1})$ is not directly measurable. In fact, this is what we are trying to estimate with language modeling. So we directly model and minimize this in stage 1 with CRT, using the same architecture as in stage 2, to ensure that our inductive biases transfer. In Figure 10 (bottom right), we plot the stage 2 losses per-position for a 211M parameter auto-regressive model between our baseline tokenizer and CRT. These tokenizers have the same codebook size, utilization, and per-position entropy. However, we see that CRT lowers loss uniformly across token positions, with most of the loss reduction happening at the end of the sequence. This implies that the CRT tokenizer has less estimated entropy  [^4] Note that the variance of the model’s final loss is very low ($\leq 10^{-3}$), so this is a significant phenomenon (Figure 8)..
@@ 5 Conclusion & Future Work
We systematically study the tradeoff between compression and generation for image tokenization and reveal several key insights: First, we demonstrate that the relationship between compression and generation quality is nuanced: smaller models actually benefit from more aggressive compression, even at the cost of reconstruction fidelity. Second, we provide a principled framework for analyzing this trade-off through the lens of scaling laws, showing consistent patterns across multiple orders of magnitude in computational budget. Finally, we introduce Causally Regularized Tokenization (CRT), a method for optimizing tokenizers specifically for autoregressive generation. We introduce a causal inductive biases during tokenizer training which substantially improves compute efficiency (2-3x) and parameter efficiency (4x reduction) compared to previous approaches. Notably, these gains are most pronounced in the resource-constrained regime, making the method particularly valuable for practical applications. Our results highlight the importance of considering the interaction between different components for multi-stage machine learning pipelines: rather than optimizing each stage independently, we achieve significant gains by baking in inductive biases in early stages. This principle suggests several promising directions for future work, for example extending to other architectures (e.g. diffusion models) or other modalities (e.g. video and audio tokenizers).
Supplementary Material
Structure of Appendix. Appendix A documents full training details for all experiments and architectures. Appendix B discusses the limitations of our experimental methodology. Appendix C provides more detailed scaling analysis across a range of experiments in the main paper. Appendix D (Figures 15-19) provides non-cherrypicked qualitative generations across a range of ImageNet categories for our best model.
@@ A.1 Stage 1: VQGAN training
We largely follow the training recipes from [9] and [43] to train our tokenizer, with modifications that we state here. The VQGAN architecture is exactly that of [9], with the following standard loss terms (not including our CRT loss):
@@
The perceptual loss comes from LPIPS [58]. The $\ell_{2}$ loss is defined as $\mathcal{L}_{2}(x,y)=\|x-y\|_{2}^{2}$ between the reconstruction and the input. The weights for each loss are: $\lambda_{VQ}=1.0$, $\lambda_{GAN}=0.5$, $\lambda_{Perceptual}=1.0$, $\lambda_{L2}=1.0$. We start $\lambda_{GAN}$ at 0. After 20k iterations, we anneal $\lambda_{GAN}$ to 0.5 using a cosine schedule for 2k iterations.
GAN Loss Details. The GAN loss uses the PatchGAN [21] as a discriminator with the Wasserstein GAN loss function from [2] directly on reconstructions from the auto-encoder to force the output to look “realistic.” This is the only “reference-free” image loss used. Using improved StyleGAN and StyleGAN2 [24] discriminators ended up with worse qualitative results in our setting.
Codebook Loss Details. Following [53, 43], we project from 256 to 8 dimensions before doing codebook lookup and perform codebook lookup with cosine distance. We use two losses, a quantization error loss and a commitment loss, both proposed in [49], with a commitment $\beta=0.25$. We do not use entropy loss, as it was unnecessary to prevent codebook collapse in our setup. We find that training with a long learning rate warmup period is crucial for scaling beyond 16k codes, otherwise performance saturates at 2.1 rFID and codebook utilization stagnates.
Optimization Parameters. We use Adam [26] with learning rate $41e-$ 4 and $\beta=(0.9,0.95)$. AdamW with weight decay $0.1$ performs identically. We use the same optimization parameters for the PatchGAN discriminator. We use with linear warmup from $51e-$ 5 for 3k iterations, batch size 128 and a constant learning rate. We found that this warmup is critical for good performance, especially for codebook length and training duration scaling. Cosine learning rate gave very similar results, and constant learning rate allows for continued training. We train for a total of 400k iterations unless stated otherwise.
@@ A.2 Stage 2: Auto-regressive training
Architecture. Following [43], we use the Llama 2 architecture [46] with QK-layer normalization [5] for stability. For some 576 tokens/image runs, we also needed to use z-loss [4, wortsman2023small] with a coefficient of $41e-$ 4 for stability. We find that this did not impact the performance of stable runs. The
Optimization. We perform data augmentation using ten-crop, allowing us to pre-tokenize the train set during stage 2 training. We use AdamW, betas $(0.9,0.95)$ and a cosine learning with schedule with a linear warmup (5k iterations) and a batch size of 256. Our max learning rate is $33e-$ 3, with a pre-warmup learning rate of $43e-$ 4, annealing our learning rate to 0. We also include a class token dropout rate of 0.1, where we replace the class-token with a randomly initialized learnable dummy token to enable classifier-free guidance. We train for a total of 375k iterations unless otherwise stated (we often sweep this parameter for stage 2 scaling law experiments).
Class-conditioning and Classifier-Free Guidance (CFG). For class-conditioning, we prepend a special class token to image token sequences at train time. At inference time, we use the relevant class token to start the sequence to generate an image of a particular class. We sample with next-token prediction without top-$k$ or top-$p$ sampling. All our results use CFG at inference time to further emphasize classifier guidance. To do this, we have two inference procedures, one with the class-token and one using the aforementioned dummy token. At each timestep, these yield logits $\ell_{u}$ for the unconditional logit, and $\ell_{c}$ for the class-conditioned logit. To get the next token, we sample from the combined logit $\ell_{cfg}=\ell_{u}+(\ell_{c}-\ell_{u})\alpha$, where $\alpha$ is known at the CFG scale. This kind of sampling was initially used for diffusion [18], but has seen success in auto-regressive image generation as well [56, 43] and is used commonly.
CFG trades off fidelity and diversity (related to precision and recall respectively) of generation, and we found that the optimal CFG scale was also pareto optimal for precision and recall. To choose our CFG scale at inference time, we split a 50k subset of the train set (not held out during training) to compute FID against. We select the optimal FID against samples from $\alpha\in\{1.5,1.75,2.0,2.25\}$ and use those generations against the validation set. We do this because CFG can have a large effect on FID score [30, 43, 45], and optimizing directly against the ImageNet validation set can pollute our evaluation. However, in separate experiments, we find that optimal CFG against this train split and the validation set are identical in practice.
@@ A.3 Scaling law sweep details
To produce scaling laws (e.g. Figure 3), we sweep training compute by changing the number of total training iterations. Because we are in the repeated data setting, unlike [23], and we are using a cosine learning rate schedule, we need to evaluate the end points of separate, independent training runs to produce these curves. The iteration counts we sweep over are $\{15k,22.5k,45k,90k,187.5k,375k,750k,1.5M,2.25M\}$ for all models. We compute estimates for training FLOPs using the approximation from [23]. The architecture parameters for our sweep can be found in Table 4.
Figure 11: Stage 2 Validation Loss by Dateset Size. Here we study the data scale bottleneck by varying the total size of the dataset. Every model was trained for 250k iterations at batch size 256. We see that our dataset is large enough to demonstrate scaling laws at the parameter range studied.
@@ Appendix B Limitations
Figure 12: Individual model scale comparison for Figures 3 and 8. Here we show how stage 2 gFID scales with respect to train FLOPs with both the CRT and baseline tokenizer. We see that with CRT tokens, stage 2 learns faster and also ends up more performant across all model sizes, especially at smaller scales (note the log-scaled y-axis).
Figure 13: Full scaling law plots for pareto curves from Figure 7. Here we show individual model scaling (similar to Figure 3) at three different stage 1 codebook sizes. 1k outperforms 131k at low compute regimes and the trend reverses at high compute regimes. Note that 16k outperforms both even though it has worse rFID that the 131k codebook size tokenizer (2.21 rFID vs 1.72 rFID).
Figure 14: Full scaling law plots for tokens per image scaling exploration (Figure 4). Here we show individual model scaling (similar to Figure 3) at three different tokens per image values. We see the same trend but in more detail, where 576 performs poorly in low compute regimes (and with small models) but outpeforms 400 and 256 tokens per image at high compute regimes. Note that 256 tokens per image is compute optimal up until the 2.5 gFID barrier is reached,at which point the tokenizer reconstruction quality becomes a bottleneck.
Figure 15: Non-cherrypicked generation examples from CRTopt-AR-775M (ImageNet 256x256). CFG scale = 2.0, class index 250, class name Siberian Husky
Figure 16: Non-cherrypicked generation examples from CRTopt-AR-775M (ImageNet 256x256). CFG scale = 2.0, class index 972, class name cliff
Figure 17: Non-cherrypicked generation examples from CRTopt-AR-775M (ImageNet 256x256). CFG scale = 2.0, class index 207, class name Golden Retriever
Figure 18: Non-cherrypicked generation examples from CRTopt-AR-775M (ImageNet 256x256). CFG scale = 2.0, class index 90, class name lorikeet
Figure 19: Non-cherrypicked generation examples from CRTopt-AR-775M (ImageNet 256x256). CFG scale = 2.0, class index 29, class name axolotl
Figure 20: Non-cherrypicked generation examples from CRTopt-AR-775M (ImageNet 256x256). CFG scale = 2.0, class index 297, class name sloth bear
Figure 21: Non-cherrypicked generation examples from CRTopt-AR-775M (ImageNet 256x256). CFG scale = 2.0, class index 248, class name husky
Figure 22: Non-cherrypicked generation examples from CRTopt-AR-775M (ImageNet 256x256). CFG scale = 2.0, class index 973, class name coral reef
Figure 23: Non-cherrypicked generation examples from CRTopt-AR-775M (ImageNet 256x256). CFG scale = 2.0, class index 993, class name gyromitra
Figure 24: Non-cherrypicked generation examples from CRTopt-AR-775M (ImageNet 256x256). CFG scale = 2.0, class index 388, class name giant panda
Figure 25: Non-cherrypicked generation examples from CRTopt-AR-775M (ImageNet 256x256). CFG scale = 2.0, class index 387, class name red panda
Figure 26: Non-cherrypicked generation examples from CRTopt-AR-775M (ImageNet 256x256). CFG scale = 2.0, class index 417, class name balloon
Figure 27: Non-cherrypicked generation examples from CRTopt-AR-775M (ImageNet 256x256). CFG scale = 2.0, class index 279, class name Arctic fox
Figure 28: Non-cherrypicked generation examples from CRTopt-AR-775M (ImageNet 256x256). CFG scale = 2.0, class index 291, class name lion
@@
Our work has a few specific limitations, largely related to the availability of compute. First, the largest model we test is 775M, which we do in order to prevent overfitting. This is close to the range of current state-of-the-art image generation models, such as [11] which ranges from 800M parameters to 8B parameters, but is smaller than modern large language models [19, 46]. Future work could include scaling parameter counts. Second, we primarily evaluate on class-to-image synthesis and not text-to-image synthesis. We do this because FID is a much more well-behaved metric on ImageNet (see inconsistent scaling in LSUN results by parameter count in Table 3). However, this means we cannot train on web-scale datasets, and thus are in the repeated data regime, deviating from [23]. In Figure 11, we study the effects of this but varying the stage 2 training set size (375k training iterations), showing that for existing model sizes we are not yet in the data bottleneck regime. Future work could include validating an in distribution metric (e.g. CLIP FID) on a large scale text-to-image dataset and then reproducing our scaling laws.
Appendix C More Detailed Scaling Plots

### v2

@@ 4 Analysis
Factors affecting validation loss. We saw in Section 3.4 that tokens per image affected compute scaling more significantly than changing the codebook size. Why is that? We note that for a stage 2 unigram model, the loss lower bound $L_{min}$ is the empirical entropy of the codebook distribution per-position, since cross-entropy can be decomposed into $H(X)+D_{KL}(q(X)|p(X))$ and $D_{KL}\to 0$ when $q\to p$. In Figure 9 (far left), we measure this directly over the ImageNet train set over various codebook sizes.In Figure 9 (center left), we see that the per position entropy converges, even as the overall codebook entropy increases log-linearly due to uniform codebook utilization. This implies that as codebook size increases, codes become more specialized by position. This has the effect of ensuring that $L_{min}$ for a unigram model does not increase, meaning that our stage 2 model can still learn even with very large codebook sizes.
[…]
What does CRT do? $L_{min}$ for an $n$-gram model is upper-bounded by this estimate, as $H(X_{i})\geq H(X_{i}|X_{<i})$. $H(X_{i}|X_{<i})$ is not directly measurable, this quantity is what we are trying to estimate with language modeling. We directly model and minimize this in stage 1 with CRT, using the same architecture as in stage 2, to ensure that our inductive biases transfer. In Figure 9 (far right), we plot the stage 2 losses per-position for a 211M parameter auto-regressive model between our baseline tokenizer and CRT. These tokenizers have the same codebook size, utilization, and per-position entropy. However, we see that CRT lowers loss uniformly across token positions, with most of the loss reduction happening at the end of the sequence. This implies that the CRT tokenizer has less estimated entropy than baseline when using a causal entropy model.
@@ 5 Conclusion & Future Work
We systematically study the tradeoff between compression and generation for image tokenization and reveal several key insights: First, we demonstrate that the relationship between compression and generation quality is nuanced: smaller models benefit from more aggressive compression, even at the cost of reconstruction fidelity. Second, we provide a principled framework for analyzing this trade-off through the lens of scaling laws, showing consistent patterns across multiple orders of magnitude in computational budget. Finally, we introduce Causally Regularized Tokenization (CRT), a method for optimizing tokenizers specifically for autoregressive generation. We add a causal inductive biases during tokenizer training which substantially improves compute efficiency (2-3x) and parameter efficiency (4x reduction) compared to previous approaches. Notably, these gains are most pronounced in the resource-constrained regime, making the method particularly valuable for practical applications. Our results highlight the importance of considering the interaction between different components for multi-stage machine learning pipelines: rather than optimizing each stage independently, we achieve significant gains by baking in inductive biases in early stages. This principle suggests several promising directions for future work, for example extending to other architectures (e.g. diffusion models) or other modalities (e.g. video and audio tokenizers).
6 Acknowledgments
The authors thank Jonathan Hayase, Matthew Wallingford, Ludwig Schmidt, Nicholas Lourie, Will Merrill, Mohammand Rastegari, Jaskirat Singh, Lili Yu, Chunting Zhou, Artidoro Pagnoni and members of the RAIVN Lab for helpful comments and discussions throughout the course of research. VR and AF acknowledge funding by NSF IIS 1652052, IIS 1703166, DARPA N66001-19-2-4031, DARPA W911NF-15-1-0543 and gifts from Allen Institute for Artificial Intelligence, Google and Apple.
Structure of Appendix. Appendix A documents full training details for all experiments and architectures. Appendix B discusses the limitations of our experimental methodology. Appendix E provides further ablations to justify our choice of $\ell_{2}$ over cross-entropy loss with CRT. Appendix F provides more detailed scaling analysis across a range of experiments in the main paper. Appendix G (Figures 15-28) provide uncurated qualitative generations across a range of ImageNet categories for our best model.
@@ A.1 Stage 1: VQGAN training
We largely follow the training recipes from [esser2020taming] and [sun2024autoregressive] to train our tokenizer, with modifications that we state here. The VQGAN architecture is exactly that of [esser2020taming], with the following standard loss terms (not including our CRT loss):
@@
The perceptual loss comes from LPIPS [zhang2018unreasonable]. The $\ell_{2}$ loss is defined as $\mathcal{L}_{2}(x,y)=\|x-y\|_{2}^{2}$ between the reconstruction and the input. The weights for each loss are: $\lambda_{VQ}=1.0$, $\lambda_{GAN}=0.5$, $\lambda_{Perceptual}=1.0$, $\lambda_{L2}=1.0$. We start $\lambda_{GAN}$ at 0. After 20k iterations, we anneal $\lambda_{GAN}$ to 0.5 using a cosine schedule for 2k iterations.
GAN Loss Details. The GAN loss uses the PatchGAN [isola2016image] as a discriminator with the Wasserstein GAN loss function from [arjovsky2017wasserstein] directly on reconstructions from the auto-encoder to force the output to look “realistic.” This is the only “reference-free” image loss used. Using improved StyleGAN and StyleGAN2 [karras2019analyzing] discriminators ended up with worse qualitative results in our setting.
Codebook Loss Details. Following [yu2021vector, sun2024autoregressive], we project from 256 to 8 dimensions before doing codebook lookup and perform codebook lookup with cosine distance. We use two losses, a quantization error loss and a commitment loss, both proposed in [van2017neural], with a commitment $\beta=0.25$. We do not use entropy loss, as it was unnecessary to prevent codebook collapse in our setup. We find that training with a long learning rate warmup period is crucial for scaling beyond 16k codes, otherwise performance saturates at 2.1 rFID and codebook utilization stagnates.
Optimization Parameters. We use Adam [kingma2014adam] with learning rate 1e41e-4 and $\beta=(0.9,0.95)$. AdamW with weight decay $0.1$ performs identically. We use the same optimization parameters for the PatchGAN discriminator. We use with linear warmup from 1e51e-5 for 3k iterations, batch size 128 and a constant learning rate. We found that this warmup is critical for good performance, especially for codebook length and training duration scaling. Cosine learning rate gave very similar results, and constant learning rate allows for continued training. We train for a total of 400k iterations unless stated otherwise.
@@ A.2 Stage 2: Auto-regressive training
Architecture. Following [sun2024autoregressive], we use the Llama 2 architecture [touvron2023llama] with QK-layer normalization [dehghani2023scaling] for stability. For some 576 tokens/image runs, we also needed to use z-loss [chen2022pali] with a coefficient of 1e41e-4 for stability. We find that this did not impact the performance of stable runs.
Optimization. We perform data augmentation using ten-crop, allowing us to pre-tokenize the train set during stage 2 training. We use AdamW, betas $(0.9,0.95)$ and a cosine learning with schedule with a linear warmup (5k iterations) and a batch size of 256. Our max learning rate is 3e33e-3, with a pre-warmup learning rate of 3e43e-4, annealing our learning rate to 0. We also include a class token dropout rate of 0.1, where we replace the class-token with a randomly initialized learnable dummy token to enable classifier-free guidance. We train for a total of 375k iterations unless otherwise stated (we often sweep this parameter for stage 2 scaling law experiments).
Class-conditioning and Classifier-Free Guidance (CFG). For class-conditioning, we prepend a special class token to image token sequences at train time. At inference time, we use the relevant class token to start the sequence to generate an image of a particular class. We sample with next-token prediction without top-$k$ or top-$p$ sampling. All our results use CFG at inference time to further emphasize classifier guidance. To do this, we have two inference procedures, one with the class-token and one using the aforementioned dummy token. At each timestep, these yield logits $\ell_{u}$ for the unconditional logit, and $\ell_{c}$ for the class-conditioned logit. To get the next token, we sample from the combined logit $\ell_{cfg}=\ell_{u}+(\ell_{c}-\ell_{u})\alpha$, where $\alpha$ is known at the CFG scale. This kind of sampling was initially used for diffusion [ho2022classifier], but has seen success in auto-regressive image generation as well [yu2023scaling, sun2024autoregressive] and is used commonly.
CFG trades off fidelity and diversity (related to precision and recall respectively) of generation, and we found that the optimal CFG scale was also pareto optimal for precision and recall. To choose our CFG scale at inference time, we split a 50k subset of the train set (not held out during training) to compute FID against. We select the optimal FID against samples from $\alpha\in\{1.5,1.75,2.0,2.25\}$ and use those generations against the validation set. We do this because CFG can have a large effect on FID score [li2024autoregressive, sun2024autoregressive, tian2024visual], and optimizing directly against the ImageNet validation set can pollute our evaluation. However, in separate experiments, we find that optimal CFG against this train split and the validation set are identical in practice.
@@ A.3 Scaling law sweep details
To produce scaling laws (e.g. Figure 4), we sweep training compute by changing the number of total training iterations. Because we are in the repeated data setting, unlike [kaplan2020scaling], and we are using a cosine learning rate schedule, we need to evaluate the end points of separate, independent training runs to produce these curves. The iteration counts we sweep over are $\{15k,22.5k,45k,90k,187.5k,375k,750k,1.5M,2.25M\}$ for all models. We compute estimates for training FLOPs using the approximation from [kaplan2020scaling]. The architecture parameters for our sweep can be found in Table 2.
A.4 Computational Cost
Each experiment can be run on one node of 8 80GB A100s. Training the tokenizer for 40 epochs takes 3 days in this setting. The largest stage 2 model (770M) trained for 375k iterations takes 3.5 days, and the cost decreases linearly with the number of parameters and training iterations.
Figure 10: Stage 2 Validation Loss by Dataset Size. Here we study the data scale bottleneck by varying the total size of the dataset. Every model was trained for 250k iterations at batch size 256. We see that our dataset is large enough to demonstrate scaling laws at the parameter range studied.
@@ Appendix B Limitations
Our work has a few specific limitations, largely related to the availability of compute and data. First, the largest model we test is 775M, which we do in order to prevent overfitting (given that ImageNet is a finite dataset). This is close to the range of current state-of-the-art image generation models, such as [esser2024scaling] which ranges from 800M parameters to 8B parameters, but is smaller than modern large language models [hoffmann2022training, touvron2023llama]. Future work could include scaling parameter counts. Second, we primarily evaluate on class-to-image synthesis and not text-to-image synthesis. We do this because FID is a much more well-behaved metric on ImageNet (see inconsistent scaling in LSUN results by parameter count in Table 1c). However, this means we cannot train on web-scale datasets, and thus are in the repeated data regime, deviating from [kaplan2020scaling]. In Figure 10, we study the effects of this but varying the stage 2 training set size (375k training iterations), showing that for existing model sizes we are not yet in the data bottleneck regime. Future work could include validating an in distribution metric (e.g. CLIP FID) on a large scale text-to-image dataset and then reproducing our scaling laws.
Appendix C Broader Impacts
Our method, as applied in this paper, makes training of image generation models more efficient and boosts overall performance. Image generation in general has broader social impacts. For example, Deepfakes can cause personal harm. Popular text-to-image models can be seen as stealing artist data for corporate gain. In its most benign form, this method can be used to accelerate research on fully legal datasets (e.g. Creative Commons images), however that may not be how it is used in the future.
Appendix D Further Baseline Tokenizer Ablations
Table 3 shows the difference between native resolution training and resolution extrapolation (the method used in the paper). Native resolution training improves performance very little compared to extrapolating to higher resolution images. This allows us to compare a fixed tokenizer at many different compression rates, removing the confounding aspect of higher resolution training.
Figure 11 shows the result of training with and without a long learning rate warmup period. This was crucial for achieving high codebook utilization and good performance without collapse for large codebooks.
Figure 11: Effects of learning rate warmup on tokenizer performance.
Table 3: Training Tokenizers at Native Resolution. Native resolution training improves performance very little compared to extrapolating to higher resolution images. This allows us to compare a fixed tokenizer at many different compression rates, removing the confound of higher resolution training.
Tokens per Image Native resolution training rFID / gFID
400 no 1.30 / 3.76
400 yes 1.21 / 3.72
576 no 0.99 / 3.61
576 yes 0.93 / 3.66
Appendix E Further CRT Ablations
Results for different causal loss application configurations are presented in Table 4.
Table 4: Comparison between $\ell_{2}$ loss and CE loss for causal regularization. The LARP loss is also compared with components relevant to this setup (CE + stochastic VQ + cosine distance + scheduled sampling [wang2024larptokenizingvideoslearned]). The setup from [wang2024larptokenizingvideoslearned] is likely more successful in conjunction with global register tokens as opposed to patch tokens (as in our case). We report average gFID across stage 2 scales 52M,  [^1] M, 211M, and 340M trained for 250k iterations each.
Causal Loss Form rFID Avg. gFID
$\ell_{2}$ before quantization (CRT) 2.36 5.51
$\ell_{2}$ after quantization 2.61 6.37
CE after quantization ($\lambda=0.03$) 2.53 6.10
LARP (CE + SVQ + scheduled sampling) 2.67 6.08
Figure 12: Effects of CRT weight $\lambda$ on VQGAN losses. In general, increasing loss weight increases losses related to perceptual quantities (e.g. $\ell_{2}$ reconstruction, discriminator loss, and perceptual LPIPS loss). It also seems to reduce codebook losses, like VQ and commitment.
Effects of decoding order on performance. As noted in prior work [esser2020taming] raster-scan is empirically the best canonical ordering for image generation. Our results corroborate this. While CRT provides improvements in these settings, it is not enough to overcome the difference in ordering performance (>>1gFID). It is possible that a more flexible base architecture (e.g. a 1D transformer based tokenizer) than that of VQGAN would allow for more flexible re-ordering, and we believe this to be an interesting direction for future work.
Appendix F More Detailed Scaling Plots
Figure 13: Individual model scale comparison for Figures 4 and 7. Here we show how stage 2 gFID scales with respect to train FLOPs with both the CRT and baseline tokenizer. We see that with CRT tokens, stage 2 learns faster and also ends up more performant across all model sizes, especially at smaller scales (note the log-scaled y-axis).